OpenStudy (mendicant_bias):

This is another simple reduction i'm not quite getting: (Problem posted below in a moment).

OpenStudy (mendicant_bias):

Simplifying from \[mv _{_{i}}=(\frac{ 1 }{ 2 }m)(2v _{i})\cos \theta _{1}+(\frac{ 1 }{ 2 }m)(2v _{i})\cos \theta _{2}\]to\[\cos \theta _{1} = \frac{ 1 }{ 2 }\]It seems so damn simple already, how do they somehow remove the other term with cosine? By the way, the two thetas are equal and opposite in value.

OpenStudy (mendicant_bias):

Your avatar is a literal avatar. Word.

OpenStudy (anonymous):

@ajprincess WHEN YOU FINISH CAN YOU PLEASE HELP SORRY FOR CAPS KEYBOARD IS STUCK

OpenStudy (ajprincess):